# Learning to Solve Integer Linear Programs with Davis-Yin Splitting

### v3

@@ Appendix D Additional Experimental Results
In Figure 6, we show the test loss and training time per epoch for all three architectures: DYS-net, CVX-net, and PertOpt-net for 10-by-10, 20-by-20, and 30-by-30 grids. In terms of MSE loss, CVX-net and DYS-net lead to comparable performance. In the second row of Figure 6, we observe the benefits of combining the three-operator splitting with JFB Fung et al. (2022); in particular, DYS-net trains much faster. Figure 7 shows some randomly selected outputs for the three architectures once fully trained.
@@
Figure 6: Comparison of of DYS-Net, cvxpylayers Agrawal et al. (2019a), PertOptNet Berthet et al. (2020), and Blackbox Backpropagation-net (BB-Net) Pogančić et al. (2019) for three different grid sizes: $10\times 10$ (first column), $20\times 20$ (second column), and $30\times 30$ (third column). The first row shows the MSE loss vs. epochs of the testing dataset. The second row shows the training time vs. epochs.

### v4

@@ C.5 Additional Training Details for Large-Scale Shortest Path Problem
C.6 Additional Model Details for Warcraft Experiment
For PertOpt-net and BBOpt-net we again use the default hyperparameter values provided by PyEPO, namely $\lambda=10$ for BBOpt-net, number of samples equal 3, $\epsilon=1$ and Gumbel noise for PertOpt-net. We set $\gamma=0.5$ for both CVX-net and DYS-net.
C.7 Hardware
@@ Appendix D Additional Experimental Results
In Figure 6, we show the test loss and training time per epoch for all three architectures: DYS-net, CVX-net, and PertOpt-net for 10-by-10, 20-by-20, and 30-by-30 grids. In terms of MSE loss, CVX-net and DYS-net lead to comparable performance. In the second row of Figure 6, we observe the benefits of combining the three-operator splitting with JFB (Fung et al., 2022); in particular, DYS-net trains much faster. Figure 7 shows some randomly selected outputs for the three architectures once fully trained.
@@
Figure 6: Comparison of of DYS-Net, cvxpylayers (Agrawal et al., 2019a), PertOptNet (Berthet et al., 2020), and Blackbox Backpropagation-net (BB-Net) (Pogančić et al., 2019) for three different grid sizes: $10\times 10$ (first column), $20\times 20$ (second column), and $30\times 30$ (third column). The first row shows the MSE loss vs. epochs of the testing dataset. The second row shows the training time vs. epochs.